OpenStudy (anonymous):

What is the arc length when Θ = 4 pi over 7 and the radius is 5 cm?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

\(\theta\) is ? what is the situation? do you have a pic? come on, man!

OpenStudy (anonymous):

thats all i have

OpenStudy (anonymous):

is it a circle and \(\theta\) is the angle of the arc subtended?

OpenStudy (anonymous):

idk thats all they gave me

OpenStudy (mrnood):

the length of an arc is simply r * angle (when angle is in radians) Your angle is 4*pi/7 your radius is 5 length = r * angle

OpenStudy (anonymous):

so 20pi/7

OpenStudy (mrnood):

Yes

OpenStudy (anonymous):

can you help me with another one

OpenStudy (mrnood):

post it separately (and I'm going pretty soon...)